Question 4 5 pts Consider the equation nº X () + X(T) - 0 with the following boundary conditions. X(0) + X(0) -0, and X(n)

n=7
Answer #1

Solution: / Consider the equation Y2x(9)+ 12x(x) = 0 200+X/102=0 Xin+X (n)=0 Let D be the differential operator de and n=77 - C&C=0 Sin (7.92-04 Now 0 and f+44 (2=0 2461+C2=0 - AC=0, C= (a) NA -24 1 (9-10) Since ctet (A)= H+ 2742 +0, for any ack,